OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4 15 15 over 2 9 11

OpenStudy (anonymous):

Draw it

OpenStudy (anonymous):



OpenStudy (anonymous):

@UditKulka

OpenStudy (anonymous):

2x-14=16 2x=16+14 2x=30 x=15

OpenStudy (anonymous):

@madrockz got it right